OpenStudy (anonymous):

solve for x to eight decimal placed if e^(2x-7)=252. Base E is a physical constant.

OpenStudy (anonymous):

\[2x-7=\ln 252\] \[x=\frac{\ln 252+7}{2}\]

jimthompson5910 (jim_thompson5910):

e^(2x-7) = 252 ln(e^(2x-7)) = ln(252) (2x-7)ln(e) = ln(252) (2x-7)*1 = ln(252) 2x-7 = ln(252) 2x = ln(252)+7 x = ( ln(252)+7 )/2 x = 6.26471454